OpenStudy (anonymous):

Evan is painting a stool that has a cylindrical seat and four rectangular prism legs. The seat has a diameter of 8 inches and a height of 2 inches. Each leg is 2 inches by 2 inches by ten inches. How much area will Evan have to paint? Use three point one four for pi. Show your work.

ganeshie8 (ganeshie8):

surface area of seat = \(\large 2\pi r^2 + 2 \pi r h - 4(l\times w)\) = \(\large 2(3.14)(4)^2 + 2(3.14)(4)(2) - 4(2\times 2)\)

ganeshie8 (ganeshie8):

simplify

ganeshie8 (ganeshie8):

you will need to find area of the legs also and add to above seat's area